40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 6 03:30:17 2015, ericq, Update, LSC, DRFPMI Progress 
    Reply  Thu Oct 8 04:48:50 2015, ericq, Update, LSC, DRFPMI Progress 
       Reply  Thu Oct 8 13:13:20 2015, Koji, Update, LSC, DRFPMI Progress 
          Reply  Thu Oct 8 14:14:50 2015, ericq, Update, LSC, DRFPMI Progress 
             Reply  Thu Oct 8 16:48:23 2015, Koji, Update, LSC, DRFPMI Progress 
             Reply  Thu Oct 8 21:35:49 2015, rana, Update, LSC, DRFPMI Progress 
                Reply  Fri Oct 9 09:22:38 2015, ericq, Update, LSC, DRFPMI Progress oct8_allRF.pdf
                   Reply  Fri Oct 9 11:24:06 2015, Jenne, Update, LSC, DRFPMI Progress 
Message ID: 11671     Entry time: Thu Oct 8 04:48:50 2015     In reply to: 11669     Reply to this: 11672
Author: ericq 
Type: Update 
Category: LSC 
Subject: DRFPMI Progress 

Progress was made. CARM was stably locked on RF only. DARM was RF only for a few moments before I typed in a wrong number...

A change was made to the LSC model's triggering section to make the DRMI hold more reliably at zero CARM offset. Namely, the POPDC signal now has its absolute value taken before the trigger matrix. Even unwhitened, it occaisionally would somehow go negative enough to break the DRMI trigger.

AUX X laser was acting up again. As before, tweaking laser current is the temporary fix.

ELOG V3.1.3-